# Soundscape Captioning using Sound Affective Quality Network and Large Language Model

### v2

@@ IV-B Experimental setup of acoustic model
Mel feature. The setting of log Mel features follows that of PANNs [21]. The 64 Mel bins are extracted by STFT with a Hamming window length of 32ms and a hop size of 10ms.
@@
Loudness feature. Loudness features are extracted directly using the ISO_532-1.exe loudness program [^1] https://standards.iso.org/iso/532/-1/ed-1/en recommended by ISO 532-1:2017 standard (Zwicker method) [22]. The input audio files are calibrated with a “.wav” file containing the calibration signal sine 1kHz 60dB, then the loudness features are calculated in 2ms as a frame. We upload the modified Python version of this code and files to the homepage.
@@
TABLE I: Mean performance of 10 runs of SoundAQnet with convolution branches of different kernel sizes on the test set.
# Kernel size $S$ub-branch Node RFS ASC AEC pleasant eventful chaotic vibrant uneventful calm annoying monotonous
3 5 7 9 {Mel; Loudness} $n$ Time (s) Acc. (%) AUC MSE
IV-C Acoustic model results and analysis
For the acoustic model, the research questions (RQs) are as follows: 1) Can Mel features collaborate with the psycho-acoustics feature loudness to improve SoundAQnet’s performance? 2) Does introducing multiscale features improve SoundAQnet’s predictions? Is the optimal timescale for fitting different dimensions of affective assessment the same? 3) How do other methods for fusing multiscale embeddings perform? 4) How does SoundAQnet perform compared to other models? 5) Does SoundAQnet capture relationships between AEs and AQs? 6) Does SoundAQnet differentiate between ASs in AQ predictions? Among them, RQ1-3 explore the effectiveness of SoundAQnet’s structure and component; RQ4 explores the performance of SoundAQnet and other models; RQ5 analyzes relationships between AEs and AQs; RQ6 applies SoundAQnet to automatic quantitative analysis of soundscapes.
1) RQ1: Can Mel features collaborate with sound-level-related loudness features to improve model performance?
Tables II and III present the performance of SoundAQnet on the ASC task, AEC task, ISO Pleasantness (ISOP) and ISO Eventfulness (ISOE) regression tasks, and emotion-related AQ regression tasks when using different acoustic features, respectively. Due to space limitations, Tables II and III show the mean results of 10 runs, without variance. When using single-class acoustic features, SoundAQnet retains only the corresponding branches, and the number of nodes in the graph-based fusion layer is reduced by half, i.e., $n=4$.
@@ IV-C Acoustic model results and analysis
The comparison of #1 and #2 in Table II shows that for the acoustic environment-related ASC and AEC, the Mel feature is more effective than the loudness feature, consistent with previous research [43]. The reason for this is easy to understand, following the notations in Section III-A, Mel features with dimension ($N$, 64) better depict acoustic representations of different frequency bands compared to the over-compressed loudness features with dimension ($N$, 1), which makes it easier for the model to learn from Mel features. Compared with Mel-based #2, the fused feature in #3 performs better in the regression of ISOP, pleasant, chaotic, calm, and annoying, as well as in classifications of AS and AE. This indicates that introducing loudness related to human perception of sound level can effectively help Mel-based SoundAQnet on partial AQ regressions and is useful for recognizing AS and AE. Since ISOP and ISOE are linear combinations of the 8D AQs and thus offer no additional insights, we will omit their results in later experiments due to space limitations.
2) RQ2: Does introducing multiscale features help Sound-AQnet capture the acoustic environment’s sound source and scene information and its resulting human-perceived AQs?
The duration of AEs may vary between a few tens of milliseconds, such as bird chirps, and several minutes, such as music. For the 30s audio clips in this paper, the experiment participants may ignore short unpleasant sounds, but they may feel annoyed when these sounds persist throughout the audio clip. For AEs caused by various sound sources, people may also need different time scales to perceive the different AQs.
Table I shows SoundAQnet with different scale features. The scale of features, i.e., the convolution receptive field size (RFS), is determined by the convolution kernel size. For SoundAQnet with a single-scale convolution kernel, such as #1 in Table I, a Mel-based (3, 3) convolution branch and a loudness-based (3, 1) convolution branch are involved, resulting in a graph-based fusion layer with two nodes.
The performance of SoundAQnet with single-scale kernel branches is shown in Table I #1-#4. When the convolution kernel size is 7, for the ASC task, the model achieves the best result, while for the AEC task, it slightly outperforms models with single-scale convolution kernels in other sizes. For the convolution branch with a kernel size of 3, 5, 7, and 9 in SoundAQnet, the RFS of each branch’s last convolution layer relative to the input acoustic features is 76, 144, 212, and 280, respectively. The corresponding audio length time is shown in Table I. Taking the branch $S_{2}$ with a kernel size of 5 as an example, if the input audio length is less than 1.44s, SoundAQnet including $S_{2}$ will not work. From #1 to #3 in Table I, when the kernel size is increased from 3 to 7, that is, the RFS of SoundAQnet is increased from 0.76s to 2.12s, SoundAQnet’s performance on ASC is correspondingly improved, but continuing to increase the kernel size does not lead to higher accuracy. That is, for ASC and AEC tasks, the RFS at 2.12s is an appropriate resolution for SoundAQnet.
In Table I #1-#4, the emotion-related 8D AQ regression tasks achieve good results, except for #1. This indicates that the length of audio clips input to SoundAQnet needs to be greater than 0.76s to effectively capture human-perceived AQs. For #2 at the 1.44s level, SoundAQnet outperforms #1 in fitting AQs such as pleasant, uneventful, and calm. For #4 at the 2.80s level, SoundAQnet outperforms #1 in fitting AQs such as eventful, chaotic, and vibrant. The results of #1-#4 suggest that, just as people may need different time scales to perceive different AQs, SoundAQnet is time-aware on human-perceived AQ regressions, which implies that the introduction of multiscale temporal features is helpful. Finally, #5, which combines branches with different time scales, performs better on ASC and AEC tasks and also well in AQ regressions. In short, with the cooperation of small and large-size convolution kernels, SoundAQnet extracts multiscale features suitable for the target tasks, and captures acoustic environment information from multiple perspectives, thereby improving the results.
3) RQ3: What are the differences between different multiscale embedding fusion methods?
The outputs from the Mel- and loudness-based branches with kernels of (3, 5, 7, 9) are denoted as (${m3}$, ${m5}$, ${m7}$, ${m9}$) and (${l3}$, ${l5}$, ${l7}$, ${l9}$), respectively. With these representations, we can obtain $F$ by fusing them, and then feed the fusion result $F$ into the residual embedding layer. Table IV presents the performance of SoundAQnet with different fusion methods. Due to space limitations, we show the mean and variance of the MSE of 8D AQ regressions as an overall metric.
@@
TABLE IV: Mean performance of SoundAQnet with different methods for fusing Mel and loudness-based sub-branches on the test set.
# Fusion Type ASC AEC AQ regression
@@
1 Addition 94.34$\pm$0.75 0.936$\pm$0.006 1.070$\pm$0.084
2 Concat 94.47$\pm$0.57 0.934$\pm$0.005 1.068$\pm$0.089
3 Hadamard 94.65$\pm$0.51 0.937$\pm$0.004 1.071$\pm$0.092
4 $A$_Q_Mel 88.85$\pm$2.96 0.865$\pm$0.008 1.059$\pm$0.083
5 $A$_Q_Loudness 94.54$\pm$0.99 0.884$\pm$0.013 1.040$\pm$0.082
6 $A$_Q_M_Q_L 94.67$\pm$0.70 0.898$\pm$0.009 1.038$\pm$0.080
7 Graph-based 95.31$\pm$0.77 0.941$\pm$0.007 1.054$\pm$0.091
The multiscale output is given as ${M}$ = ((((${m3}$, ${m5}$, ${m7}$, ${m9}$) for Mel branches and ${L}$ = ((((${l3}$, ${l5}$, ${l7}$, ${l9}$) for loudness branches. For #1 in Table IV, ${F}$ = ${M}+{L}$. For #2, they are concatenated, ${F}$ = $Concat({M},{L}$). For #3, the Hadamard product is used, ${F}$ = ${M}$ $\odot$ ${L}$, where $\odot$ is the element-wise product. Overall, SoundAQnet performs similarly based on the fusion methods of #1-#3. Table IV #4-#6 adopt the scaled dot-product attention ($A$), a key component in the widely used Transformer [44].
@@
where $\mathbf{V}$$=$$\mathbf{K}$, and $d_{k}$ is $\mathbf{K}$’s dimension. Its essence is using the similarity between $\mathbf{Q}$ and $\mathbf{K}$ to adjust the information in $\mathbf{V}$, so a more informative $\mathbf{V}$ will lead to better results. In #4, ${M}$ acts as $\mathbf{Q}$ and ${L}$ acts as $\mathbf{K}$, then ${F}=A(M,L)$, denoted as $A$_Q_Mel, that is, Mel-based representations are used as a query to adjust loudness-based representations, and the output result mainly relies on ${L}$. The operation of #5 is the opposite of #4, and #5 performs better than #4 on ASC and AEC tasks. The reason is similar to the Mel-only and loudness-only models in Table II. Notably, both #5 and #6, which use loudness as $\mathbf{Q}$ to modulate Mel-based representations, perform better in AQ regressions. #6, which concatenates $A(M,L)$ and $A(L,M)$ in #4 and #5, shows the best result for AQ regressions. Overall, the graph-based multiscale embedding fusion SoundAQnet improves ASC and AEC performance, and shows competitive overall performance in regressions of human-perceived AQs. For all source code, see the homepage.
4) RQ4: How does the performance of SoundAQnet compare with other sound recognition models?
There are no other models similar to SoundAQnet for simultaneously modelling the AS, AE, and emotion-related affective quality (AQ). Previous studies on AQ in soundscapes often use traditional linear regression to predict some AQ response values, while recent deep-learning neural-network-based studies only focus on a few specific affective qualities [16][45]. After careful consideration, we compare SoundAQnet with deep-learning models that perform well for auditory scene and event analysis, i.e., ASC and AEC tasks, as shown in Table V.
TABLE V: Comparison of different models on the test set.
# Model Param. ASC AEC AQ regression
(M) Acc. (%) AUC MSE Mean
1 AD_CNN [27] 0.52 89.63$\pm$2.21 0.84$\pm$0.02 1.128$\pm$0.077
2 Baseline_CNN 1.01 87.87$\pm$1.76 0.92$\pm$0.01 1.315$\pm$0.144
3 Hierarchical_CNN 1.01 89.82$\pm$2.75 0.89$\pm$0.02 1.293$\pm$0.198
4 MobileNetV2 [46] 2.26 89.67$\pm$0.88 0.92$\pm$0.01 1.145$\pm$0.112
5 YAMNet [47] 3.21 88.84$\pm$1.59 0.90$\pm$0.01 1.199$\pm$0.109
6 CNN-Transformer 12.29 92.80$\pm$0.59 0.93$\pm$0.01 1.339$\pm$0.134
7 PANNs [21] 79.73 93.57$\pm$1.18 0.90$\pm$0.02 1.156$\pm$0.107
8 SoundAQnet 2.70 95.31$\pm$0.77 0.94$\pm$0.01 1.054$\pm$0.091
In Table V, #1 refers to the CNN used in the ARAUS paper [27]. CNN in #2 is the baseline for benchmarking the multiscale convolution-based SoundAQnet. It consists of 4 convolutional layers, each with 16, 32, 64, and 128 filters, and their corresponding kernel sizes of 3, 5, 7, and 9, respectively. After the convolutional layers, there are parallel ASC and AEC layers and regression layers for AQs. Hierarchical CNN (HiCNN) in #3 aims to identify AS based on the predictions of AE, exploiting the implicit hierarchical relationship between AS and AEs [28]. HiCNN modifies the ASC layer in the Baseline CNN by feeding the output of the AEC layer into the ASC. Hence, there is almost no difference in the number of parameters (Param.) between HiCNN and Baseline CNN. Compared with #2, the ASC result in #3 is better, but the AEC result is affected by the direct hierarchical connection. MobileNetV2 in #4 is a well-known lightweight CNN that uses depthwise separable convolution to reduce the computational cost [46]. YAMNet in #5 is a CNN-based baseline for AEC from Google. Given the excellent performance of Transformer-based models on audio tasks [28], #6 proposes CNN-Transformer, an encoder from Transformer [44] is added after the convolutional layer in Baseline CNN, to combine the spatial feature extraction capability of CNN with the temporal modelling capability of Transformer. Compared with #2, the introduction of Transformer encoder in #6 enhances the model’s ability to discriminate acoustic scenes and events, and improves its classification results, but its overall result on 8D AQ regressions is not as good as those of the pure CNN in #4. The reason may be that, compared with Transformer encoder modelling AQs from the hidden layer features with the global perspective, CNN relies on a fixed-size convolutional kernel and performs better in learning the hidden layer features from different local perspectives, which is beneficial for modelling the unique representation of each AQ.
Overall, the proposed SoundAQnet, which simultaneously models AS, AE, and human-perceived AQ, achieves the best results in ASC, AEC, and affect-related regression tasks with a similar number of parameters as MobileNetV2. Note that we modify the output layer of these classic models to enable them to model the acoustic environment and corresponding AQs simultaneously. For more details, please see the homepage.
@@
5) RQ5: Does SoundAQnet capture the relationship between detected audio events (AEs) and affective qualities (AQs)?
People respond affectively to various sounds in their daily environment, regardless of their nature [48]. From RQ1-4, it can be seen that SoundAQnet performs well in identifying ASs and AEs, as well as predicting the values of perceived AQs. Does SoundAQnet learn the implicit relationship between various AEs and AQs? To this end, Fig. 6 (a) shows the statistical significance of the predictions of SoundAQnet on the test set with 3576 30-second binaural audio clips to analyze the relationship between AEs and the AQs they evoke. The Shapiro-Wilk test shows that the distributions of 15 AEs and 8D AQs are non-normal ($\alpha$ $>$ $0.05$). Thus, we use Spearman’s rho for correlation analysis between AEs and AQs. The statistical results in Fig. 6 (a) show that there are significant correlations between AEs and AQs. Specifically, some AEs like ‘Bird’, ‘Animal’, ‘Outside, rural or natural’ (Outside) and ‘Silence’ have significant positive correlations with pleasantness and calm. In addition, some AEs like ‘Human sounds’, ‘Music’, and ‘Speech’ have significant positive correlations with eventful and vibrant, while some AEs, including ‘Sound of things’ and ‘Vehicle’, can significantly evoke annoyingness (Anno.) and Chaotic. This indicates SoundAQnet’s capability to capture the correlation between AEs and different AQs.

### v3

@@ V-B Experimental Setup of Acoustic Model
Mel feature. The setting of log Mel features follows that of PANNs [12]. The 64 Mel bins are extracted by STFT with a Hamming window length of 32ms and a hop size of 10ms, resulting in Mel features having 3000 frames.
@@
Loudness feature. Loudness features are extracted directly using the ISO_532-1.exe loudness program [^1] https://standards.iso.org/iso/532/-1/ed-1/en recommended by ISO 532-1:2017 standard (Zwicker method) [20]. The input audio files are calibrated with a “.wav” file containing the calibration signal, which is a sine wave at 1 kHz 60 dB. Then, the loudness features are calculated in frames of 2ms, resulting in Zwicker-loudness with 15000 frames. We upload the modified Python code and files to the homepage.
@@
V-C Comparisons to Other Models
TABLE I: Comparison of different models on the test set (batch size=32).
# Model Param Inference ASC AEC AQ regression
(M) time(ms) GPU(GB) Acc.(%) AUC MSE Mean
1 AD_CNN [32] 0.52 5.6 1.79 89.63 0.84 1.128
2 Baseline_CNN 1.01 4.3 1.18 87.87 0.92 1.315
3 Hierarchical_CNN 1.01 4.6 1.18 89.82 0.89 1.293
4 MobileNetV2 2.26 5.5 1.94 89.67 0.92 1.145
5 YAMNet 3.21 4.9 2.16 88.84 0.90 1.199
6 CNN-Transformer 12.29 4.4 1.24 92.80 0.93 1.339
7 PANNs [12] 79.73 19.0 5.67 93.57 0.90 1.156
8 SoundAQnet 2.70 18.9 3.22 95.31 0.94 1.054
There are no other models similar to SoundAQnet for simultaneously modelling the AS, AE, and emotion-related AQ. Previous studies on AQ in soundscapes often use traditional linear regression to predict some AQ response values, while recent deep-learning-based studies only focus on a few specific AQs [7][27][29]. Therefore, we compare SoundAQnet with deep-learning models that perform well for auditory scene and event analysis, i.e., ASC and AEC tasks, as shown in Table I.
In Table I, #1 refers to the CNN used in the ARAUS paper [32]. CNN in #2 is the baseline for benchmarking the multiscale convolution-based SoundAQnet. It consists of 4 convolutional layers, each with 16, 32, 64, and 128 filters, and their corresponding kernel sizes of 3, 5, 7, and 9, respectively. After the convolutional layers, there are parallel ASC and AEC layers and regression layers for AQs. Hierarchical CNN in #3 aims to identify AS based on the predictions of AE, exploiting the implicit hierarchical relationship between AS and AEs [15]. Compared with #2, the ASC in #3 is better, but the AEC is affected by the hierarchical relationship. MobileNetV2 in #4 is a lightweight CNN that uses depthwise separable convolution to reduce the computational cost [50]. YAMNet in #5 is a CNN-based baseline for AEC from Google. Given the excellent performance of Transformers on audio tasks [15], #6 proposes CNN-Transformer, an encoder from Transformer [51] is added after the convolutional layer in Baseline CNN, to combine the spatial feature extraction capability of CNN with the temporal modelling capability of Transformer. Compared with #2, the introduction of Transformer encoder in #6 enhances the model’s ability to discriminate acoustic scenes and events, and improves its classification results, but its overall result on 8D AQ regressions is not as good as those of the pure CNN in #4. The reason may be that, compared with Transformer encoder modelling AQs from the hidden layer features with the global perspective, CNN relies on a fixed-size convolutional kernel and performs better in learning the hidden layer features from different local perspectives, which is beneficial for modelling unique representations of each AQ.
Overall, the proposed SoundAQnet, which simultaneously models AS, AE, and human-perceived AQ, achieves the best results in ASC, AEC, and affect-related regression tasks with a similar number of parameters as MobileNetV2. More information on computational efficiency, such as training/inference speed, GPU, and hardware requirements, see the homepage.
V-D Ablation Studies of SoundAQnet
1) Ablation study on acoustic features
Tables II and III present the performance of SoundAQnet on ASC, AEC, ISO Pleasantness (ISOP) and ISO Eventfulness (ISOE) regression, and emotion-related AQ regression tasks when using different acoustic features, respectively. They are the mean results over 10 runs. When using single-class acoustic features, SoundAQnet retains only the corresponding branches, and the number of nodes in the graph-based fusion layer is reduced by half, i.e., $n=4$.
@@ V-D Ablation Studies of SoundAQnet
The comparison of #1 and #2 in Table II shows that for the acoustic environment-related ASC and AEC, the Mel feature is more effective than the loudness feature, consistent with previous research [31]. The reason for this is easy to understand, following the notations in Section IV-A, Mel features with dimension ($N$, 64) include spectral information, which is essential for recognizing sounds, while this essential information is lost in over-compressed loudness features with dimension ($N$, 1). Compared with Mel-based #2, the fused feature in #3 performs better in the regression of ISOP, pleasant, chaotic, calm, and annoying, as well as in classifications of AS and AE. This indicates that introducing Zwicker-loudness explicitly can help Mel-based SoundAQnet on partial AQ regressions, and AS and AE classification. The effect of adding loudness is stronger for ISOP (the valence axis in Fig. 1), which is in line with expectations [2][20]. Since ISOP and ISOE are linear combinations of the 8D AQs and thus do not offer additional insight, we will omit their results in the following experiments due to space limitations.
2) Ablation study on multiscale sub-branches
The duration of AEs may vary between a few tens of milliseconds, such as bird chirps, and several minutes, such as music. AQs along the valence dimension, ISOP, may be determined by short sounds that are identified as calming or annoying. AQs that are aligned with high arousal, high ISOE, such as eventful, chaotic, and vibrant, are related to multiple changes in the sound environment and require longer time windows to be identified. This ablation study aims to explore whether these expected relationships can be linked to the kernel sizes used in different branches.
@@
TABLE IV: Mean performance of 10 runs of SoundAQnet with convolution branches of different kernel sizes on the test set.
# Kernel size SSub-branch Node RFS ASC AEC pleasant eventful chaotic vibrant uneventful calm annoying monotonous
3 5 7 9 {Mel; Loudness} $n$ Time (s) Acc. (%) AUC MSE
Table IV shows the performance of SoundAQnet using features with different scales. The scale of features, i.e., the convolution receptive field size (RFS), is determined by the convolution kernel size. The performance of SoundAQnet with single-scale kernel branches is shown in #1-#4 of Table IV. For the convolution branch with a kernel size of 3, 5, 7, and 9, the RFS of each branch’s last convolution layer relative to the input acoustic features is 76, 144, 212, and 280, respectively. The corresponding RFS in seconds is shown in Table IV. From #1 to #3, when the kernel size is increased from 3 to 7, that is, the RFS is increased from 0.76s to 2.12s, SoundAQnet’s performance on ASC and AEC is improved, but continuing to increase the kernel size does not lead to higher accuracy.
In Table IV #1-#4, the emotion-related 8D AQ regression tasks achieve good results, except for #1. This indicates that the length of audio clips input to SoundAQnet needs to be greater than 0.76s to effectively capture human-perceived AQs. For #2 at the 1.44s RFS, SoundAQnet outperforms #1 in predicting AQs pleasant, uneventful, calm, and annoying, but a small decrease in performance is seen when increasing the kernel size further. For #4 at the 2.80s RFS, SoundAQnet outperforms options with smaller kernel sizes in predicting AQs eventful, chaotic, and vibrant. The results of #1-#4 suggest that SoundAQnet is time-window-aware, just as people may need different time scales to perceive different AQs. Finally, #5, which combines branches with different kernel sizes and RFS, performs better on ASC and AEC tasks while still performing reasonably well in most AQ regressions. Other combinations of $S_{1}$, $S_{2}$, $S_{3}$, and $S_{4}$ have been explored, with more details shown in the homepage. In short, with the cooperation of small and large-size convolution kernels, SoundAQnet extracts multiscale features suitable for the target tasks, and captures acoustic environment information from multiple perspectives, thereby improving the results.
TABLE V: Mean performance of SoundAQnet with different methods for fusing Mel and loudness-based sub-branches on the test set.
# Fusion: ($\mathbf{M}$}, $\mathbf{L}$}) ASC AEC AQ regression
@@
[…]
3) Ablation study on multiscale embedding fusion
The multiscale output is given as $\mathbf{M}$} = ((m3{m3}, m5{m5}, m7{m7}, m9{m9}) for Mel branches and $\mathbf{L}$} = ((l3{l3}, l5{l5}, l7{l7}, l9{l9}) for loudness branches. We can obtain the fusion result $F$ by fusing these outputs, and then feed $\mathbf{F}$} into the residual embedding layer. Table V shows SoundAQnet’s performance with different fusion methods. The mean and variance of the MSE of 8D AQ regressions are shown as an overall metric. For #1 in Table V, they are added; for #2, they are concatenated; for #3, the Hadamard product is used, where $\odot$ is the element-wise product. SoundAQnet performs similarly based on the fusion of #1-#3. #4-#6 adopt the scaled dot-product attention ($Att$), a key component in the widely used Transformer [51].
@@
where $\mathbf{V}$$\mathbf{K}$, and $d_{k}$ is $\mathbf{K}$’s dimension. In this case, the similarity between $\mathbf{Q}$ and $\mathbf{K}$ is exploited to adjust the information in $\mathbf{V}$, so a more informative $\mathbf{V}$ can lead to better results. In #4, $\mathbf{M}$} acts as $\mathbf{Q}$ and $\mathbf{L}$} acts as $\mathbf{K}$, Mel-based representations are used as a query to adjust loudness-based representations, and the output result mainly relies on $\mathbf{L}$}. The operation of #5 is the opposite of #4, and #5 performs better than #4 on ASC and AEC tasks. The reason is similar to the Mel-only and loudness-only models in Table II. Notably, both #5 and #6, which use loudness as $\mathbf{Q}$ to modulate Mel-based representations, perform better in AQ regressions. #6, which concatenates $Att()$ in #4 and #5, shows the best result for AQ regressions. #7 uses the attention feature fusion (AFF) for variable-length audio in contrastive language-audio pretraining (CLAP) [52].
$X_{fusion}=\alpha X_{global}+(1-\alpha)X_{local}$ (7)
where $\alpha=f_{AFF}(X_{global},X_{local})$ is a factor obtained by AFF [52]. Unlike CLAP, which only uses Mel features, SoundAQnet uses two types of features. Thus, for CLAP-AFF-based SoundAQnet, Mel-based branches and loudness-based branches calculate the corresponding $X_{fusion\_Mel}$ and $X_{fusion\_loudness}$, respectively. Then, these two are concatenated and fed into a 1-layer multilayer perceptron (MLP) for fusion. Overall, the graph-based multiscale embedding fusion improves ASC and AEC performance, and shows competitive overall performance in regressions of human-perceived AQs. Source code for these fusions can be found in the homepage.
4) The impact of network depth
SoundAQnet contains three layers of Conv2D blocks, each referring to VGG [38] and consisting of two convolution layers. Table VI presents the number of parameters (Param.), multiply-accumulate operations (MACs), and GPU memory required to train SoundAQnet with different depths.
TABLE VI: SoundAQnet with different numbers of layers. (Batch size=32)
# blocks # layers Para. MACs GPU ASC AEC AQ regression
(M) (G) (GB) Acc. (%) AUC MSE Mean
2 4 1.53 19.58 11.67 93.46 0.926 1.092
3 6 2.70 27.68 13.54 95.31 0.941 1.054
4 8 7.36 93.02 21.48 95.60 0.938 1.058
5 10 25.92 667.97 58.69 95.44 0.933 1.060
In Table VI, as SoundAQnet’s depth increases, its computational overhead increases significantly but does not bring better results. SoundAQnet with the default number of blocks, three, balances model complexity, computational overhead, and achieves competitive performance in ASC, AEC, and AQ regression tasks. Due to space limitations, please also refer to the homepage for code and results for different dilation rates.
@@
V-E Correlation Study between AEs and AQs with SoundAQnet
People respond affectively and rate AQs based on recognition of various AEs, which are more related to AQ along the arousal axis [53]. This section answers the question whether SoundAQnet performs well in predicting AQs by implicitly learning the relationship between AEs and AQs. To investigate, Fig. 6 (a) shows the statistical significance of the predictions given by SoundAQnet on the test set of 3576 30-second binaural audio clips to analyze the relationship between AEs and the AQs they evoke. The Shapiro-Wilk test shows that the distributions of 15 AEs and 8D AQs are non-normal ($\alpha$ >> $0.05$). Thus, we use Spearman’s rho for correlation analysis between AEs and AQs. The statistical results in Fig. 6 (a) show that there are significant correlations between AEs and AQs. Specifically, some AEs like ‘Bird’, ‘Animal’, ‘Outside, rural or natural’ (Outside) and ‘Silence’ have significant positive correlations with pleasantness and calm. In addition, some AEs like ‘Human sounds’, ‘Music’, and ‘Speech’ have significant positive correlations with eventful and vibrant, while some AEs, including ‘Sound of things’ and ‘Vehicle’, can significantly evoke annoyingness (Anno.) and Chaotic. This indicates SoundAQnet’s capability to capture the correlation between AEs and different AQs.